OpenStudy (anonymous):

Find the prime factorization of 50. A. 1 x 2 x 5 x 5 B. 2 x 5 x 5 C. 2 x 25 D. 2 x 2 x 5

OpenStudy (anonymous):

B

OpenStudy (anonymous):

50 / 2 = 25 25 / 5 = 5 5 / 5 = 1 the second column are the prime factors